A 135 kg scaffold is 6.20 m long. It is hanging with two wires, one from each end. A 560 kg box sits 1.60 m from the left end. What is the tension in the left hand side wire?(g = 9.8 m/s2)
Answer #1

considering torque about right end

Mg ( L/2) + mg ( L - d) = T * L

135* 9.8 * 3.1 + 560* 9.8* ( 6.2 - 1.6) = T * 6.2

T = 4733.24 N

=======

Comment before rate in case any doubt, will reply for sure.. goodluck